# On Spheres with $k$ Points Inside

### v2

@@ Abstract
We generalize a classical result by Boris Delaunay that introduced Delaunay triangulations. In particular, we prove that for a locally finite and coarsely dense generic point set $A$ in ${{\mathbb{R}}}^{d}$, every generic point of ${{\mathbb{R}}}^{d}$ belongs to exactly $\binom{d+k}{d}$ simplices whose vertices belong to $A$ and whose circumspheres enclose exactly $k$ points of $A$. We extend this result to the cases in which the points are weighted, and when $A$ contains only finitely many points in ${{\mathbb{R}}}^{d}$ or in ${{\mathbb{S}}}^{d}$. Furthermore, we use the result to give a new geometric proof for the fact that volumes of hypersimplices are Eulerian numbers.
@@
Triangulations, higher-order Delaunay triangulations, hypertriangulations, Delone sets, $k$-sets, Worpitzky’s identity, hypersimplices.
@@ 1 Introduction
In the seminal paper [4], Boris Delaunay (also spelled Delone) introduced the Delaunay triangulation of a finite point sets using simplices with empty circumspheres. His construction can be reformulated as follows: for a (finite and generic) point set, $A\subseteq{{\mathbb{R}}}^{d}$, the simplices with vertices in $A$ that contain no points of $A$ inside their circumspheres cover the convex hull of $A$ in one layer. In this paper, we generalize Delaunay’s construction and prove similar properties for simplices with circumspheres that enclose exactly $k$ points of $A$, for some fixed non-negative integer $k$. We call these simplices the $k$-hefty simplices of $A$.
We introduce the main concepts we will work with. A set $A\subseteq{{\mathbb{R}}}^{d}$ is locally finite if every closed ball contains at most a finite number of the points of $A$, and it is coarsely dense if every closed half-space contains at least one and therefore infinitely many of the points of $A$. If $A$ has both properties, we call it a thin Delone set; compare with the more restrictive class of Delone sets, which are uniformly discrete and relatively dense, meaning the smallest inter-point distance is bounded away from 00, and the radius of the largest empty ball is bounded away from $\infty$. We call $A$ generic if no $d+1$ of its points lie on a common hyperplane, and no $d+2$ of its points lie on a common $(d-1)$-sphere. Any $(d-1)$-sphere bounds a closed $d$-ball and thus partitions ${{\mathbb{R}}}^{d}$ into points inside the sphere (in the interior of the ball), points on the sphere, and points outside the sphere (in the complement of the ball). Assuming $A$ is generic, there is a unique $(d-1)$-sphere passing through any $d+1$ points of $A$, which we call the circumscribed sphere of the $d$-simplex spanned by the points.
@@
Let $k$ be a non-negative integer and $A\subseteq{{\mathbb{R}}}^{d}$ a generic thin Delone set or a generic finite set. A $d$-simplex with vertices in $A$ is $k$-hefty if exactly $k$ points of $A$ lie inside the circumsphere of the $d$-simplex.
For example, the 00-hefty simplices are the top-dimensional simplices in the Delaunay triangulation of $A$, and $k$-hefty simplices with $k>0$ are related to the cells in higher-order Delaunay triangulations [2, 6]. Our main result is Theorem 2.3, which we restate here in less technical terms:
@@
Let $k$ be a non-negative integer and $A\subseteq{{\mathbb{R}}}^{d}$ a generic thin Delone set. Then the $k$-hefty simplices of $A$ cover ${{\mathbb{R}}}^{d}$ in exactly $\binom{d+k}{d}$ layers.
We also prove versions of this theorem for finitely many points, points on the $d$-dimensional sphere, and weighted points. In addition, we apply the covering multiplicities to get a new proof that the volumes of hypersimplices are Eulerian numbers, and to get new proofs for some bounds on $k$-sets.
The paper is organized as follows. In Section 2, we introduce the main definitions, prove the main result for thin Delone sets (Theorem 2.3) and finite sets (Corollary 2.5), and formulate their local versions (Theorem 2.7). In Section 3, we apply the results to obtain a new proof for the fact that volumes of hypersimplices are Eulerian numbers and new proofs for old bounds on $k$-sets. In the concluding Section 4, we discuss extensions of the results to points in hyperbolic and spherical spaces and to points with real weights in Euclidean space.
@@ 2 Hefty Simplices in Euclidean Space
This section presents the main result of this paper, which is stated for infinite and finite point sets in Euclidean space. We begin with the main technical lemma before stating and proving the main theorem.
2.1 Main Technical Lemma
For technical reasons we first show that the $k$-hefty simplices of a thin Delone set $A$ are “locally uniform” in size. Specifically, we prove an upper bound for the radii of spheres that enclose a fixed point, $x\in{{\mathbb{R}}}^{d}$, as well as at most $k$ points of $A$. To this end, we write $B(x,R)$ for the closed ball with center $x$ and radius $R$, and note that the number of points of $A$ in this ball goes to infinity when $R$ goes to infinity.
@@ 2.1 Main Technical Lemma
Let $A\subseteq{{\mathbb{R}}}^{d}$ be coarsely dense, $k$ a non-negative integer, and $x\in{{\mathbb{R}}}^{d}$. Then there exists $R=R(x,A,k)$ such that if $x$ is inside a sphere that is not fully contained in $B(x,R)$, then there are at least $k+1$ point of $A\cap B(x,R)$ inside this sphere.
@@
Without loss of generality, assume $x=0$. For every unit vector, $u\in{{\mathbb{S}}}^{d-1}$, the open halfspace of points $y$ that satisfy $(y,u)>0$ contains infinitely many points of $A$. It follows that the function $f_{u}\colon(0,\infty)\to{{\mathbb{Z}}}$ that maps $r>0$ to the number of points of $A$ inside the sphere with center $ru$ and radius $r$ is non-decreasing and unbounded.
We introduce $g\colon{{\mathbb{S}}}^{d-1}\to{{\mathbb{R}}}$ defined by $g(u)=\inf\{r>0\mid f_{u}(r)\geq k+1\}$ and claim that $g$ is bounded. To derive a contradiction, suppose $g$ is unbounded, and let $u_{1},u_{2},\ldots$ be a sequence of unit vectors with $g(u_{n})\geq n$. Since ${{\mathbb{S}}}^{d-1}$ is compact, there is a subsequence that converges to a vector $u_{0}\in{{\mathbb{S}}}^{d-1}$. Let $S$ be the sphere with radius $g(u_{0})+1$ and center $(g(u_{0})+1)u_{0}$. By construction, there are at least $k+1$ points of $A$ inside $S$. Since these points are (strictly) inside the sphere, there is a sufficiently small ${\varepsilon}>0$ such that moving the center of the sphere by at most ${\varepsilon}$ while adjusting its radius so the origin remains on the sphere, retains at least $k+1$ point of $A$ inside the sphere. But this contradicts the unboundedness of $g$ as there are points $u_{i}$ in the subsequence that are within distance ${\varepsilon}$ from $u_{0}$ with $g(u_{i})$ much larger than $g(u_{0})+1$.
@@
As an immediate consequence of Lemma 2.1, the circumsphere of any $k$-hefty simplex that encloses $x$ is completely contained in $B(x,R)$.
@@ 2.2 Global Covering
Our first goal is to generalize the classic result of Delaunay that the 00-hefty simplices of every thin Delone set cover ${{\mathbb{R}}}^{d}$ in one layer; that is: every point of ${{\mathbb{R}}}^{d}$ is contained in at least one 00-hefty simplex and almost every point of ${{\mathbb{R}}}^{d}$ is contained in exactly one 00-hefty simplex. Specifically, we show that for every generic thin Delone set, $A\subseteq{{\mathbb{R}}}^{d}$, the family of $k$-hefty simplices covers ${{\mathbb{R}}}^{d}$ $\binom{d+k}{d}$ times. We call $\binom{d+k}{d}$ the $k$-th covering number and note that it depends on the dimension, $d$, and the parameter, $k$, but not on the set $A$. We call $x\notin A$ generic with respect to $A$ if $A\cup\{x\}$ is generic. Almost every point $x\in{{\mathbb{R}}}^{d}$ is generic with respect to a generic thin Delone set, $A$. To see this, observe that by local finiteness of $A$ there are only countably many hyperplanes spanned by $d$ points each or spheres spanned by $d+1$ points each, so the union of these hyperplanes and spheres has Lebesgue measure zero.
@@
Step 1. Letting $k$ be a non-negative integer and $A\subseteq{{\mathbb{R}}}^{d}$ a generic thin Delone set, we prove that here is a constant $c=c(k,A)$ such that any point that is generic with respect to $A$ is contained in exactly $c$ $k$-hefty simplices of $A$. Write ${{cover}_{k}{({x,A})}}$ for the number of $k$-hefty simplices of $A$ that contain $x$. By Lemma 2.1, ${{cover}_{k}{({x,A})}}$ is finite. Indeed, every $k$-hefty simplex of $A$ that contains $x$ must select its vertices from the finitely many points inside the ball $B(x,2M)$. To show that ${{cover}_{k}{({x,A})}}$ is the same for all generic points, we move $x$ continuously from one point to another. The only time ${{cover}_{k}{({x,A})}}$ can change is when $x$ passes through the boundary of a $k$-hefty simplex. It suffices to show that for every $(d-1)$-simplex, $\Delta$, with vertices in $A$, the number of $k$-hefty simplices with facet $\Delta$ is the same on both sides of $\Delta$.
Consider the line, $L$, that consists of all points equidistant to the vertices of $\Delta$ and mark each point $y\in L$ with the number of points of $A$ inside the sphere with center $y$ that passes through the vertices of $\Delta$. This partitions $L$ into labeled intervals, and since $A$ is generic, the labels of two consecutive intervals differ by exactly one. Fix a left to right direction on $L$, move $y$ in this direction, and observe that the portion of space inside the sphere centered at $y$ that lies to the left of the hyperplane spanned by $\Delta$ shrinks, while the portion to the right of this hyperplane grows. The transitions from an interval labeled $k+1$ to another labeled $k$ are in bijection with the $k$-hefty simplices with facet $\Delta$ to the left of $\Delta$. Indeed, as $y$ makes the transition, there is a point of $A$ that passes from inside to outside the sphere centered at $y$, so this point is to the left of $\Delta$. Similarly, the transitions from an interval labeled $k$ to another labeled $k+1$ are in bijection with the $k$-hefty simplices with facet $\Delta$ to the right of $\Delta$. There are equally many transitions of either kind because the labels go to infinity on both sides. This proves that ${{cover}_{k}{({x,A})}}$ does not depend on $x$.
@@
Figure 1: Two circles in the $1$-parameter family of circles that pass through the endpoint of the edge $\Delta$. Both are the circumcircles of $k$-hefty triangles, with $k=3$ in the shown case. As we move the center from left to right, every point that leaves the inside of the circle lies to the left of $\Delta$, and every point that enters the inside of the circle lies to the right of $\Delta$.
Step 2. We strengthen by showing that the constant in Step 1 depends on $d$ and $k$ but not on $A$. Specifically, we prove that for every dimension $d$ and non-negative integer $k$, there exists a number $f(d,k)$ such that for any generic thin Delone set, $A\subseteq{{\mathbb{R}}}^{d}$, and any point, $x\in{{\mathbb{R}}}^{d}$, that is generic with respect to $A$, belongs to exactly $f(d,k)$ $k$-hefty simplices of $A$.
It suffices to show that for two thin Delone sets, $A$ and $A^{\prime}$, and two points, $x,x^{\prime}\in{{\mathbb{R}}}^{d}$, that are generic with repect to both sets, ${{cover}_{k}{({x,A})}}={{cover}_{k}{({x^{\prime},A^{\prime}})}}$. By Lemma 2.1, there exists $R>0$ such that if a sphere encloses $x$ and a point outside $B(x,R)$, then there are at least $k+1$ points of $A\cap B(x,R)$ inside this sphere. It follows that the circumspheres of all $k$-hefty simplices that enclose $x$ are contained in $B(x,R)$. Similarly, let $R^{\prime}$ be the constant from Lemma 2.1 for $x^{\prime}$ and $A^{\prime}$. We construct a new thin Delone set, $A^{\prime\prime}$: picking points $y$ and $y^{\prime}$ at distance larger than $R+R^{\prime}$ from each other, we let $A^{\prime\prime}\cap B(y,R)$ be a translate of $A\cap B(x,R)$ and $A^{\prime\prime}\cap B(y^{\prime},R^{\prime})$ a translate of $A^{\prime}\cap B(x^{\prime},R^{\prime})$. We perturb the points if necessary to achieve genericity, and add more points outside $B(y,R)$ and $B(y^{\prime},R^{\prime})$ until $A^{\prime\prime}$ is thin Delone. By choice of $R$ and $R^{\prime}$ we have ${{cover}_{k}{({x,A})}}={{cover}_{k}{({y,A^{\prime\prime}})}}$, because every $k$-hefty simplex of $A$ whose circumsphere encloses $x$ translates into a $k$-hefty simplex of $A^{\prime\prime}$ whose circumsphere encloses $y$, and vice versa. Similarly, we have ${{cover}_{k}{({x^{\prime},A^{\prime}})}}={{cover}_{k}{({y^{\prime},A^{\prime\prime}})}}$. Finally, ${{cover}_{k}{({x,A})}}={{cover}_{k}{({x^{\prime},A^{\prime}})}}$ because ${{cover}_{k}{({y,A^{\prime\prime}})}}={{cover}_{k}{({y^{\prime},A^{\prime\prime}})}}$ as proved in Step 1.
@@
Figure 2: Before perturbation, the points of $A$ lie on three half-lines emanating from the origin. We show six $2$-hefty triangles, and emphasize two of them by shading.
Step 3. We provide an explicit example of a generic thin Delone set, $A\subseteq{{\mathbb{R}}}^{d}$, and a point, $x\in{{\mathbb{R}}}^{d}$, with ${{cover}_{k}{({x,A})}}=\binom{d+k}{d}$. Specifically, we prove that for every dimension $d$ and non-negative integer $k$, there exists a generic thin Delone set $A\subseteq{{\mathbb{R}}}^{d}$, and a point $x$ generic with respect to $A$, such that ${{cover}_{k}{({x,A})}}=\binom{d+k}{d}$.
Let $S$ be a regular $d$-simplex with vertices $v_{0},v_{1},\ldots,v_{d}$ and barycenter $0=\tfrac{1}{d+1}\sum_{i}v_{i}$, and $A$ the set of points of the form $iv_{j}$, for integers $i\geq 1$ and $j=0,1,\ldots,d$. We call $A$ a radial set and the points with fixed $j$ a direction of $A$; see Figure 2. Set $x=0$. Every $k$-hefty simplex of $A$ that contains 00 has exactly one vertex from each direction. Let $i_{0}v_{0},i_{1}v_{1},\ldots,i_{d}v_{d}$ be the vertices of such a simplex. The number of points of $A$ inside its circumsphere is $\sum_{j}(i_{j}-1)$, so $\sum_{j}i_{j}=k+d+1$. Enumerating these simplices is the same as writing $d+k+1$ as an ordered sum of $d+1$ positive integers, and there are exactly $\binom{d+k}{d}$ ways to do that. To complete the proof, we perturb $A$ while making sure that $d+1$ points span a simplex that contains 00 before the perturbation iff they span such a simplex after the perturbation. This completes the proof of our main result.
2.3 The Finite Case
For a finite set, $A\subseteq{{\mathbb{R}}}^{d}$, the covering multiplicity of Theorem 2.3 holds sufficiently deep inside the set but acts only as an upper bound near the fringes. To formalize these claims, we introduce a parametrized generalization of the convex hull: for each integer $k\geq 0$, the $k$-hull of $A$, denoted ${{H}_{k}{({A})}}$, is the common intersection of all closed half-spaces that miss at most $k$ of the points in $A$. Clearly, ${{H}_{0}{({A})}}={{\rm conv\,}{A}}$, and ${{H}_{k}{({A})}}\supseteq{{H}_{k+1}{({A})}}$ for every $k$. By the Centerpoint Theorem of discrete geometry [5, Section 4.1], ${{H}_{k}{({A})}}\neq\emptyset$ if $k<\frac{n}{d+1}$, in which $n={{\#}{A}}$ is the number of points in $A$.

### v3

@@ Abstract
We generalize the classic definition of Delaunay triangulation and prove that for a locally finite and coarsely dense generic point set, $A\subseteq{{\mathbb{R}}}^{d}$, the $d$-simplices whose vertices belong to $A$ and whose circumscribed spheres enclose exactly $k$ points of $A$ cover ${{\mathbb{R}}}^{d}$ exactly $\binom{d+k}{d}$ times. Similarly, the subset of such simplices incident to a point in $A$ cover any small enough neighborhood of that point exactly $\binom{d+k-1}{d-1}$ times. We extend this result to the cases in which the points are weighted and when $A$ contains only finitely many points in ${{\mathbb{R}}}^{d}$ or in ${{\mathbb{S}}}^{d}$. Using these results, we give new proofs of classic results on $k$-facets, old and new combinatorial results for hyperplane arrangements, and a new proof for the fact that the volumes of hypersimplices are Eulerian numbers.
@@
Triangulations, higher-order Delaunay triangulations, hypertriangulations, Delone sets, $k$-sets, Worpitzky’s identity, hypersimplices, $k$-facets, hyperplane arrangements.
@@ 1 Introduction
In the seminal paper [8], Boris Delaunay (also spelled Delone) introduced the Delaunay triangulation of a finite point sets using simplices with empty circumscribed spheres. His construction can be reformulated as follows: for a (finite and generic) point set, $A\subseteq{{\mathbb{R}}}^{d}$, the simplices with vertices in $A$ that contain no points of $A$ inside their circumscribed spheres cover the convex hull of $A$ in one layer. In this paper, [^1] An earlier version of this paper appeared in [10], which this version extends by generalizing the main theorem to weighted points in Section 2.5 and to points and weighted points on the sphere in Sections 3.1 and 3.2, reproving a result by Clarkson about the maximum number of local minima in the $k$-level of a hyperplane arrangement in Section 4.2, and giving a tight bound on the maximum number of minimum helf chambers in a half-space arrangement in Section 4.3.we generalize Delaunay’s construction and prove similar properties for simplices with circumscribed spheres that enclose exactly $k$ points of $A$, for some fixed non-negative integer $k$. We call these simplices the $k$-hefty simplices of $A$.
We begin by introducing the main concepts. A set $A\subseteq{{\mathbb{R}}}^{d}$ is locally finite if every closed ball contains at most a finite number of the points of $A$, and it is coarsely dense if every closed half-space contains at least one and therefore infinitely many points of $A$. If $A$ has both properties, we call it a thin Delone set; compare with the more restrictive class of Delone sets, which are uniformly discrete and relatively dense, meaning the smallest inter-point distance is bounded away from 0, and the radius of the largest empty ball is bounded away from $\infty$. We call $A$ generic if no $d+1$ of its points lie on a common hyperplane, and no $d+2$ of its points lie on a common (d1)(d-1)-sphere. Any (d1)(d-1)-sphere bounds a closed $d$-ball and thus partitions ${{\mathbb{R}}}^{d}$ into points inside the sphere (in the interior of the ball), points on the sphere, and points outside the sphere (in the complement of the closed ball). Assuming $A$ is generic, there is a unique (d1)(d-1)-sphere passing through any $d+1$ points of $A$, which we call the circumscribed sphere of the $d$-simplex spanned by the points.
@@
Let $k$ be a non-negative integer and $A\subseteq{{\mathbb{R}}}^{d}$ a generic thin Delone set or a generic finite set. A $d$-simplex with vertices in $A$ is $k$-hefty if exactly $k$ points of $A$ lie inside the circumscribed sphere of the $d$-simplex.
For example, the 0-hefty simplices are the top-dimensional simplices in the Delaunay triangulation of $A$, and $k$-hefty simplices with $k>0$ are related to the cells in higher-order Delaunay triangulations [5, 11]. Our main results are Theorems 2.3 and 2.5, which we restate here in less technical terms:
@@
Let $k$ be a non-negative integer, $A\subseteq{{\mathbb{R}}}^{d}$ a generic thin Delone set, and $a\in A$. Then the $k$-hefty simplices of $A$ cover ${{\mathbb{R}}}^{d}$ exactly $\binom{d+k}{d}$ times, and the $k$-hefty simplices incident to $a$ cover any sufficiently small neighborhood of $a$ $\binom{d+k-1}{d-1}$ times.
More specifically, almost every point of ${{\mathbb{R}}}^{d}$ is covered by exactly $\binom{d+k}{d}$ $k$-hefty simplices, while boundary points of $k$-hefty simplices are of course contained in more than this number of such simplices. Similarly, almost every point sufficiently close to $a\in A$ is covered by exactly $\binom{d+k-1}{d-1}$ $k$-hefty simplices with vertex $a$, while boundary point are again contained in more than this number of such simplices. We also prove versions of this theorem for finitely many points, weighted points, and points on the $d$-dimensional sphere. In addition, we apply the results to get new proofs of old and new results on $k$-facets, arrangements of hyperplanes, arrangements of hemispheres, and the volume of hypersimplices.
The paper is organized as follows. Section 2 introduces the main definitions and proves the main result for thin Delone sets (Theorems 2.3 and 2.5), which it then extends to finite sets (Theorem 2.7) and weighted points (Theorem 2.11). Section 3 considers hefty simplices on the $d$-dimensional sphere, and generalizes the local and global covering results for points (Theorem 3.2) and for weighted points (Theorem 3.5). Section 4 applies the results on covering to give new proofs of old and new results in discrete geometry: a new proof of Alon and Győry’s bound on the number of at-most-$k$-sets (Proposition 4.1), a new proof of a sharpened version of Lovász Lemma (Proposition 4.3), a new proof of Clarkson’s result on the number of minima in a hyperplane arrangement (Proposition 4.5), a proof of a new bound on the maximum number of minimum heft cells in a half-space arrangement (Theorem 4.7), and a new proof of the fact that the volumes of hypersimplices are Eulerian numbers (Theorem 4.10). Section 5 offers concluding remarks.
@@ 2 Hefty Simplices in Euclidean Space
This section presents the main result of this paper, which we state for infinite sets and then extend to finite sets and sets of weighted points in Euclidean space. We begin with the main technical lemma before stating and proving the main theorem.
2.1 First a Technical Lemma
For technical reasons we first show that the $k$-hefty simplices of a thin Delone set $A$ are “locally uniform” in size. Specifically, we prove an upper bound for the radii of spheres that enclose a fixed point, $x\in{{\mathbb{R}}}^{d}$, as well as at most $k$ points of $A$. To this end, we write $B(x,R)$ for the closed ball with center $x$ and radius $R$, and note that the number of points of $A$ in this ball goes to infinity when $x$ is fixed and $R$ goes to infinity.
@@ 2.1 First a Technical Lemma
Let $A\subseteq{{\mathbb{R}}}^{d}$ be coarsely dense, $k$ a non-negative integer, and $x\in{{\mathbb{R}}}^{d}$. Then there exists $R=R(x,A,k)$ such that if $x$ is inside a sphere that is not fully contained in $B(x,R)$, then there are at least $k+1$ points of $A\cap B(x,R)$ inside this sphere.
@@
Without loss of generality, assume $x=0$. For every unit vector, $u\in{{\mathbb{S}}}^{d-1}$, the open half-space of points $y$ that satisfy $(y,u)>0$ contains infinitely many points of $A$. It follows that the function $f_{u}\colon(0,\infty)\to{{\mathbb{Z}}}$ that maps $r>0$ to the number of points of $A$ inside the sphere with center $ru$ and radius $r$ is non-decreasing and unbounded.
We introduce $g\colon{{\mathbb{S}}}^{d-1}\to{{\mathbb{R}}}$ defined by $g(u)=\inf\{r>0\mid f_{u}(r)\geq k+1\}$ and claim that $g$ is bounded. To derive a contradiction, suppose $g$ is unbounded, and let $u_{1},u_{2},\ldots$ be an infinite sequence of unit vectors with $g(u_{n})\geq n$. Since ${{\mathbb{S}}}^{d-1}$ is compact, there is a subsequence that converges to a vector $u_{0}\in{{\mathbb{S}}}^{d-1}$. Let $S$ be the sphere with radius $g(u_{0})+1$ and center $(g(u_{0})+1)u_{0}$. By construction, there are at least $k+1$ points of $A$ inside $S$. Since these points are (strictly) inside the sphere, there is a sufficiently small ${\varepsilon}>0$ such that moving the center of the sphere by at most $\varepsilon$} while adjusting its radius so the origin remains on the sphere, retains at least $k+1$ points of $A$ inside the sphere. But this contradicts the unboundedness of $g$ as there are points $u_{i}$ in the subsequence that are within distance $\varepsilon$} from $u_{0}$ with $g(u_{i})$ much larger than $g(u_{0})+1$.
@@
As an immediate consequence of Lemma 2.1, the circumscribed sphere of any $k$-hefty simplex of $A$ that encloses $x$ is completely contained in $B(x,R)$, in which $R$ depends on $x$, AA, and $k$.
@@ 2.2 Global Covering
Our first goal is to generalize the classic result of Delaunay, which asserts that the 0-hefty simplices of a generic thin Delone set cover ${{\mathbb{R}}}^{d}$ in one layer; that is: every point of ${{\mathbb{R}}}^{d}$ is contained in at least one 0-hefty simplex and almost every point of ${{\mathbb{R}}}^{d}$ is contained in exactly one 0-hefty simplex. Specifically, we show that for every generic thin Delone set, $A\subseteq{{\mathbb{R}}}^{d}$, the $k$-hefty simplices cover ${{\mathbb{R}}}^{d}$ $\binom{d+k}{d}$ times. We call $\binom{d+k}{d}$ the $k$-th covering number and note that it depends on the dimension, $d$, and the parameter, $k$, but not on the set $A$. We call $x\notin A$ generic with respect to $A$ if $A\cup\{x\}$ is generic. For a given generic thin Delone set $A$, almost every point $x\in{{\mathbb{R}}}^{d}$ is generic with respect to $A$. To see this, observe that by local finiteness of $A$ there are only countably many hyperplanes and spheres spanned by $d$ and $d+1$ points, respectively, so the union of these hyperplanes and spheres has Lebesgue measure zero.
@@
Step 1. Letting $k$ be a non-negative integer and $A\subseteq{{\mathbb{R}}}^{d}$ a generic thin Delone set, we prove that there is a constant such that any point generic with respect to $A$ is contained in exactly this constant number of $k$-hefty simplices of $A$. Write ${c}_{k}{({x,A})}$} for the number of $k$-hefty simplices of $A$ that contain $x$. By Lemma 2.1, ${c}_{k}{({x,A})}$} is finite. Indeed, every $k$-hefty simplex of $A$ that contains $x$ must select its vertices from the finitely many points inside the ball $B(x,2M)$. To show that ${c}_{k}{({x,A})}$} is the same for all generic points, we move $x$ continuously from one point to another. The only time ${c}_{k}{({x,A})}$} can change is when $x$ passes through the boundary of a $k$-hefty simplex. It suffices to show that for every (d1)(d-1)-simplex, $\Delta$, with vertices in $A$, the number of $k$-hefty simplices with facet $\Delta$ is the same on both sides of $\Delta$.
Consider the line, $L$, that consists of all points equidistant to the vertices of $\Delta$ and mark each point $y\in L$ with the number of points of $A$ inside the sphere with center $y$ that passes through the vertices of $\Delta$. This partitions $L$ into labeled intervals, and since $A$ is generic, the labels of two consecutive intervals differ by exactly one. Fix a left to right direction on $L$, move $y$ in this direction, and observe that the portion of space inside the sphere centered at $y$ that lies to the left of the hyperplane spanned by $\Delta$ shrinks, while the portion to the right of this hyperplane grows. The transitions from an interval labeled $k+1$ to another labeled $k$ are in bijection with the $k$-hefty simplices with facet $\Delta$ to the left of $\Delta$. Indeed, as $y$ makes the transition, there is a point of $A$ that passes from inside to outside the sphere centered at $y$, so this point is to the left of $\Delta$. Similarly, the transitions from an interval labeled $k$ to another labeled $k+1$ are in bijection with the $k$-hefty simplices with facet $\Delta$ to the right of $\Delta$. There are equally many transitions of either kind because the labels go to infinity on both sides. This proves that ${c}_{k}{({x,A})}$} does not depend on $x$.
@@
Figure 1: Two circles in the 11-parameter family of circles that pass through the endpoint of the edge $\Delta$. Both are the circumcircles of $k$-hefty triangles, with $k=3$ in the case displayed. As we move the center from left to right, every point that leaves the inside of the circle lies to the left of $\Delta$, and every point that enters the inside of the circle lies to the right of $\Delta$.
Step 2. We strengthen the claim proved in Step 1 by showing that the constant depends on $d$ and $k$ but not on $A$. Specifically, we prove that for every dimension, $d$, and non-negative integer, $k$, there exists a number ${c}_{k}{({d})}$} such that for any generic thin Delone set, $A\subseteq{{\mathbb{R}}}^{d}$, any point $x\in{{\mathbb{R}}}^{d}$ generic with respect to $A$ belongs to exactly ${c}_{k}{({d})}$} $k$-hefty simplices of $A$.
It suffices to show that for two thin Delone sets, $A$ and $A^{\prime}$, and two points, $x,x^{\prime}\in{{\mathbb{R}}}^{d}$, that are generic with respect to both sets, ${{c}_{k}{({x,A})}}={{c}_{k}{({x^{\prime},A^{\prime}})}}$. By Lemma 2.1, there exists $R>0$ such that if a sphere encloses $x$ and a point outside $B(x,R)$, then there are at least $k+1$ points of $A\cap B(x,R)$ inside this sphere. It follows that the circumscribed spheres of all $k$-hefty simplices that enclose $x$ are contained in $B(x,R)$. Similarly, let $R^{\prime}$ be the constant from Lemma 2.1 for $x^{\prime}$ and $A^{\prime}$. We construct a new thin Delone set, $A^{\prime\prime}$, by picking points $y$ and $y^{\prime}$ at distance larger than $R+R^{\prime}$ from each other, and letting $A^{\prime\prime}\cap B(y,R)$ and $A^{\prime\prime}\cap B(y^{\prime},R^{\prime})$ be translates of $A\cap B(x,R)$ and $A^{\prime}\cap B(x^{\prime},R^{\prime})$, respectively. We perturb the points if necessary to achieve genericity, and add more points outside $B(y,R)$ and $B(y^{\prime},R^{\prime})$ until $A^{\prime\prime}$ is thin Delone. By choice of $R$ and $R^{\prime}$ we have ${{c}_{k}{({x,A})}}={{c}_{k}{({y,A^{\prime\prime}})}}$, because every $k$-hefty simplex of $A$ whose circumscribed sphere encloses $x$ translates into a $k$-hefty simplex of $A^{\prime\prime}$ whose circumscribed sphere encloses $y$, and vice versa. Similarly, we have ${{c}_{k}{({x^{\prime},A^{\prime}})}}={{c}_{k}{({y^{\prime},A^{\prime\prime}})}}$. Finally, ${{c}_{k}{({x,A})}}={{c}_{k}{({x^{\prime},A^{\prime}})}}$ because ${{c}_{k}{({y,A^{\prime\prime}})}}={{c}_{k}{({y^{\prime},A^{\prime\prime}})}}$ as proved in Step 1.
@@
Figure 2: Before perturbation, the points of $A$ lie on three half-lines emanating from the origin. We show all six 22-hefty triangles, and emphasize two of them by shading.
Step 3. We provide an explicit example of a generic thin Delone set, $A\subseteq{{\mathbb{R}}}^{d}$, and a point, $x\in{{\mathbb{R}}}^{d}$, with ${{c}_{k}{({d})}}=\binom{d+k}{d}$. Specifically, we prove that for every dimension $d$ and non-negative integer $k$, there exists a generic thin Delone set $A\subseteq{{\mathbb{R}}}^{d}$, and a point $x$ generic with respect to $A$, such that ${{c}_{k}{({x,A})}}=\binom{d+k}{d}$.
Consider a regular $d$-simplex with vertices $v_{0},v_{1},\ldots,v_{d}$ and barycenter $0=\tfrac{1}{d+1}\sum_{j}v_{j}$, and $A$ the set of points of the form $iv_{j}$, for integers $i\geq 1$ and $j=0,1,\ldots,d$. We call $A$ a radial set and the points with fixed $j$ a direction; see Figure 2. Set $x=0$. Every $k$-hefty simplex of $A$ that contains 0 has exactly one vertex in each direction. Letting $i_{0}v_{0},i_{1}v_{1},\ldots,i_{d}v_{d}$ be these vertices, the number of points inside the circumscribed sphere is $\sum_{j}(i_{j}-1)=k$. Enumerating these simplices is the same as writing $k$ as an ordered sum of $d+1$ non-negative integers, and there are exactly $\binom{d+k}{d}$ ways to do that. To complete the proof, we perturb $A$ while making sure that $d+1$ points span a simplex that contains 0 before the perturbation iff they span such a simplex after the perturbation. This completes the proof of the first claim in our Main Theorem.
2.3 Local Covering
By Theorem 2.3, the $k$-hefty simplices of a thin Delone set cover ${{\mathbb{R}}}^{d}$ without gap an integer number of times. However, beyond one dimension, it is generally not possible to split such a cover into sub-covers (separate layers) that enjoy the same property. As an example, consider Figure 3, which shows one point surrounded by five others in ${{\mathbb{R}}}^{2}$. Assuming the six points are generic and the remaining points of the thin Delone set—which are not shown—are sufficiently far from the six, the point at the center is incident to five $1$-hefty triangles—which are shown. Globally, the $1$-hefty triangles cover any small neighborhood of this point three times, but it is not possible to decompose this cover into three; see [17] for indecomposable coverings in other contexts. It is however possible to split it locally into two covers. The simplices that prevent a local split into three separate layers are all incident to a common vertex, and we prove that the hefty simplices incident to a common point cover any small neighborhood a fixed number of times. This is the second claim in our Main Theorem.
[…]
Figure 3: Six points of a Delone set, in which one is surrounded by five others, and the remaining points lie outside the circumcircles defined by these points. The five $1$-hefty triangles incident to the surrounded point cover any small neighborhood of that point twice, but they cannot be split into two disjoint covers of such a neighborhood.